OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −x − 6 y = −x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (swissgirl):

ok if its parallel all it means is that it has the same slope 3x-y=-3 can be rewritten as y=3x+3 which is in the form of y=mx+b where m is our slope so what is the slope for this line?

OpenStudy (campbell_st):

the easy way is the substitute the point into the 4 lines... only one will return the correct answer or take you line and rewrite it in point slope form y = 3x + 3 find the slope